OpenStudy (anonymous):

Need help urgently; it's financial math! Will fan and medal anyone who can help!

OpenStudy (anonymous):

When she turned 18, Kayla purchased 110 shares of Stock A for $38 per share; 88 shares of Stock B for $109 per share; 78 shares of Stock C for $57 per share; and 18 shares of Stock D for $98 per share. At age 36, Kayla sold Stock A for $5,019, Stock B for $11,490, Stock C for $6,131, and Stock D for $1,219 with annual dividends of $45 per year. Which stock provided Kayla with the greatest return on her investment? Stock A Stock B Stock C Stock D

OpenStudy (whpalmer4):

Let's look at them one by one. Stock A was bought for $38/share (110 shares) and sold for $5019. How much did stock A cost to purchase?

OpenStudy (anonymous):

38 * 110 - 4,180. So $4,180 to purchase Stock A?

OpenStudy (whpalmer4):

Yes, that's correct. If it cost $4,180 to purchase, and Kayla received $5,019 when selling, how much was the return on the investment in stock A?

OpenStudy (anonymous):

$839

OpenStudy (whpalmer4):

I believe you are correct :-) How about stock B? 88 shares @ $109, sold for $11,490

OpenStudy (anonymous):

Stock B: 88 * 109 = $592 Stock C: 78 * 57 = 4,446 Stock D: 18 * 98 = 1,764

OpenStudy (anonymous):

For the costs

OpenStudy (anonymous):

So if she sold Stock B for $11,490 - $9,592 = $ 1,898

OpenStudy (anonymous):

Stock C: $6,131 - 4,446 = $1,685

OpenStudy (anonymous):

@whpalmer4 I'm not sure how to calculate Stock D because it has annual dividends of $45 per year

OpenStudy (anonymous):

$1,219 - 1,764 = -545 which makes her in the negatives so she sold it for less than she invested but she also get the $45 annual dividends per year

OpenStudy (whpalmer4):

okay, so, she got $45 annually in dividends. How many years?

OpenStudy (anonymous):

It doesn't say but it did mention earlier that she is 36 when she sells but she had bought at age 18. So should I use that?

OpenStudy (anonymous):

I don't think it would justify the cost though because 18*45 would only make her $810

OpenStudy (anonymous):

I think she makes the most off of Stock B

OpenStudy (whpalmer4):

Sorry, wrote numbers down in wrong order! basis on stock D was 18*98 = 1764 pocketed 810 in dividends sold for 1219 net: 1219+810-1764 = 265 Yes, B looks like the best investment. Easy to pick the good investments afterwards, it's before that is tricky :-)

OpenStudy (anonymous):

@whpalmer4 thank you so much for all of your help and for helping me to understand it! Thank you! Fanned and medal!